40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Sep 4 19:39:36 2014, Tara, Evan, DailyProgress, FSS, TTFSS OLTFs pdhTFs.pdf
    Reply  Sun Sep 7 11:54:21 2014, Evan, DailyProgress, FSS, TTFSS OLTFs, residual frequency noise pdhTFs.pdfoltfNorth.pdfoltfSouth.pdffrnoise.pdfnoiseBudget.pdf
       Reply  Mon Sep 8 04:20:41 2014, rana, DailyProgress, FSS, TTFSS OLTFs, residual frequency noise 
       Reply  Wed Oct 2 19:09:31 2019, anchal, DailyProgress, FSS, TTFSS OLTFs; Repeated this 5 year old measurement FSS_OLTF.pdf
          Reply  Thu Jan 9 19:00:19 2020, anchal, DailyProgress, FSS, TTFSS OLTFs FSS_OLTF.pdf
             Reply  Wed Jan 22 14:16:28 2020, anchal, DailyProgress, FSS, TTFSS OLTFs FSS_OLTF.pdf
                Reply  Wed Jan 22 16:00:26 2020, anchal, DailyProgress, FSS, TTFSS OLTFs with Maximum Gain FSS_OLTF.pdf
                   Reply  Thu Jan 23 19:23:03 2020, anchal, DailyProgress, FSS, NFSS Boost Stage NorthFSSBosstStageTF.pdf
Message ID: 1502     Entry time: Thu Sep 4 19:39:36 2014     Reply to this: 1504
Author: Tara, Evan 
Type: DailyProgress 
Category: FSS 
Subject: TTFSS OLTFs 

With the cavities locked, Tara and I took OLTFs of the PDH loops.

Below 100 kHz, we used the SR785 with a 70.7 mVpk excitation. Above 30 kHz, we used the HP4395A with a 22.4 mVrms excitation (these are the "HF" traces on the attached plot).

Before taking these TFs, we turned the loop gains up as high as possible without making the loops saturate.

  • For north, the gains were 900 common and 900 fast, and the incident power was 1.26(5) mW.
  • For south, the gains were 632 common and 770 fast, and the incident power was 1.05(2) mW.

We injected the excitation on common EXC. We then measured the TF which takes common OUT2 to common OUT1. This is almost the OLTF of the loop, except for an AD829 between OUT2 and OUT1 which has a gain of −4 V/V.

Currently I'm not sure how to explain the magnitude discrepancy between the SR785 and HP4395 measurements. Both OUT1 and OUT2 have a 50 Ω output impedance, so I would expect the impedance difference between the SR785 and HP4395 would cancel out in this measurement.

Attachment 1: pdhTFs.pdf  127 kB  Uploaded Thu Sep 4 20:41:58 2014  | Hide | Hide all
pdhTFs.pdf
ELOG V3.1.3-